OpenStudy (anonymous):

Given 3x^2-15x+ (-5) at which point will the curve have slope (a) 0 (b) 2 (c) - 2 All the answers MUST be points. (a) The point where the slope is 0 is . (b) The point where the slope is 2 is . (c) The point where the slope is - 2 is .

OpenStudy (anonymous):

is this calculus? i guess take derivative=slope

OpenStudy (anonymous):

a) d/dx(3x^2-15x-5)=0 b) d/dx(3x^2-15x-5)=2 c) d/dx(3x^2-15x-5)=-2

OpenStudy (anonymous):

then for each case you will get a value of x which you can plug into original eqn to get the y